### v2

@@ Abstract
Over the past decade, visual gaze estimation has garnered growing attention within the research community, thanks to its wide-ranging application scenarios. While existing estimation approaches have achieved remarkable success in enhancing prediction accuracy, they primarily infer gaze directions from single-image signals and discard the huge potentials of the currently dominant text guidance. Notably, visual-language collaboration has been extensively explored across a range of visual tasks, such as image synthesis and manipulation, leveraging the remarkable transferability of large-scale Contrastive Language-Image Pre-training (CLIP) model. Nevertheless, existing gaze estimation approaches ignore the rich semantic cues conveyed by linguistic signals and priors in CLIP feature space, thereby yielding performance setbacks. In pursuit of making up this gap, we delve deeply into the text-eye collaboration protocol and introduce a novel gaze estimation framework in this paper, referred to as GazeCLIP. Specifically, we intricately design a linguistic description generator to produce text signals with coarse directional cues. Additionally, a CLIP-based backbone that excels in characterizing text-eye pairs for gaze estimation is presented. This is followed by the implementation of a fine-grained multi-modal fusion module aimed at modeling the interrelationships between heterogeneous inputs. Extensive experiments on three challenging datasets demonstrate the superiority of the proposed GazeCLIP which surpasses the previous approaches and achieves the state-of-the-art estimation accuracy.
keywords:
@@ 1 Introduction
During the past decade, visual gaze estimation has garnered ever-increasing attention thanks to its widespread application scenarios, such as saliency detection [3, 4], virtual reality[5], human-robot interaction[40, 41], medical diagnosis[42], driver fatigue estimation[6, 7] and other areas[1, 2]. Gaze estimation hammers at inferring accurate gaze direction of a human face or eyes from the input image. While the practical potential of this visual task is undeniably promising, it nonetheless grapples with substantial challenges. For example, the use of diverse camera views inevitably introduces unpredictable variations in lighting conditions, colors, and facial appearances, resulting in severe performance degradation. To struggle with this issue, several existing studies[17][23, 24, 25][51, 52, 53, 54] have made lots of efforts to marry the stability and robustness to regression algorithms. An additional catalyst for the thriving of investigations on gaze estimation is the increasing availability of large-scale datasets [12, 13, 14, 15, 16] which consist of varied subjects including lighting conditions, head poses and so on.
@@
Inspired by the huge success of deep learning in computer vision[44, 45, 46, 47], an increasing number of researchers are shifting their focus towards modeling complex mappings between input signals and gaze labels[16, 17][22]. Nonetheless, these methods typically operate on single-modal facial images as input, making them susceptible to the challenging task of extracting gaze-specific features amidst the presence of unrelated regions (e.g., the nose and mouth) scattered throughout the input image as different regions are prone to confuse the model on what to prioritize in learning. A general model for video-based remote POG estimation[8] evades this barrier by mining eye geometric characteristics. Albeit the impressive progress, these conventional geometry-modelling approaches exhibit several drawbacks, such as limited laboratory environment and poor generalization performance to new subject, thereby leading to a surge in the booming of appearance-based gaze estimation algorithms, harnessing the capabilities of cutting-edge deep learning models. Specifically, to enhance the estimation performance, recent learning-based protocols [9, 10, 11] excel in extracting useful features from raw facial or eye images through the design of diverse high-capacity CNN models, see Figure 1(a). Furthermore, the inescapable drastic variations introduced by diverse camera views and individual characteristics (e.g. gender, skin color and glasses) [12][22] easily get the model trapped into failure of learning highly-nonlinear relationships, resulting in overfitting and subpar performance. Therefore, there is a large room for further exploration of gaze regression model with enhanced learning capabilities.
@@
Recently, a cohort of large-scale pre-traiend linguistic-vision models [31, 32, 33, 34] have come to dominate the forefront of artificial intelligence, and trigger the vogue of multi-modal learning. Leveraging copious unlabeled data in conjunction with a contrastive learning framework helps yield robust characteristics. Particularly, Contrastive Language-Image Pre-Training (CLIP) [18] elicits the foremost attention within the research community, and has been adopted by widespread downstream vision tasks, such as image generation and visual question answering [48, 49, 50]. Specifically, CLIP’s backbone is comprised of set of transformer blocks, and is trained using extensive image-text pairs (400 million). A well-trained CLIP empowers the model with fascinating capabilities for modeling inter-relationships between language and visual modalities. The primary advantage of incorporating CLIP lies in its implicit enrichment of features with the nuanced semantics embedded in natural language. For example, object detection [19], semantic segmentation [20], and image classification [21] have achieved notable success by harnessing the guidance ability of the prominent CLIP model.
@@
Despite the remarkable progress attained by previous problems, it is noteworthy that the field of gaze estimation neglects the substantial benefits of leveraging language guidance. Indeed, the task of regressing raw face/eye images into continuous gaze vectors presents a intractable challenge. Therefore, a "divide-and-conquer" idea naturally emerges. It tries to first predict several coarse-grained directions of gaze and then utilize the coarse-granularity pseudo labels to guide the fine-grained gaze estimation (accurate direction vector). Moreover, consider the potential for the gaze estimator to yield less satisfactory results when the target gaze direction deviates from the camera’s central axis, as opposed to cases where the gaze is directly aimed at the camera, we believe that the estimation accuracy can be notably enhanced through the guidance of coarse-grained pseudo directions. Motivated by the booming of CLIP-equipped approaches in other vision tasks, we are pioneering an unprecedented attempt to leverage the collaborative effects of text-image semantic cues for gaze estimation. More specifically, coarse directions (pseudo labels) can be readily represented using natural language and subsequently encoded through pre-trained CLIP encoders, thereby enhancing the discriminative embeddings with robust guidance ability.
@@
While linguistic prompts offer effective practice for expressing high-level and coarse-grained directions, devising appropriate prompts from facial/eye images can pose a significant challenge, as articulating the nuances between input images in words can often be a difficult undertaking. Unlike image classification tasks, the generation of meaningful language descriptions (such as gaze direction prompt) using the full-fledged BLIP [43] is inherently impractical as the training of BLIP fully ignores the samples of gaze faces, thereby impeding the applicability of CLIP in the field of gaze estimation.
@@
To address the aforementioned challenges, this paper delves into the intricate process of formulating suitable prompts and proceeds to carefully design a language-vision protocol (GazeCLIP). This protocol aims at enhancing the semantic space of the gaze estimation model. Specifically, we pre-define the text prompt as “A photo of a face gazing [class]”, where [class] represents a set of base direction(front, up, down, left, right). In order to make the constructed prompt conform to the language prior knowledge of the CLIP model, we directly use the zero-shot CLIP model to assign a direction for each image. We calculate the similarity between each input image and the sentence composed of the five directions mentioned above, and take the direction with the highest similarity as the text input. On top of the generated prompt-image pairs, we presence a cross-attention-based module to align the representations from two heterogeneous modalities. Following the fine-grained fusion of multi-modal features through transformer blocks and cross-attention condensers, the resulting refined image embedding are combined with the original image embedding together to act as the input of final regression head. As depicted in Fig. 1, Unprecedentedly, our emphasis lies on harnessing the guidance provided by textual signals, rather than solely struggling to design image encoders.
@@
We evaluate our proposed model on three challenging public dataset benchmarks for gaze estimation. Extensive experiments demonstrate that GazeCLIP achieves state-of-the-art performance in all three datasets. Particularly, compared with existing approaches, the GazeCLIP gains significant improvement with an angular error decrease of 0.5° on average(4.1° → 3.6°, 7.7° → 7.3°, 5.3° → 4.7°). In a nutshell, the overall contributions of our GazeCLIP are threefold:
@@
  • In Use. A novel text-guided gaze estimation method is proposed in this paper, namely GazeCLIP, to enhance the performance of gaze estimation, which is inspired by strong generalization capability of CLIP model in various downstream vision tasks. To the best of our knowledge, this is the first effort to distil rich knowledge from full-fledged pre-trained language-vision model for guiding the learning of gaze estimation network.
@@
  • SOTA Performance. Extensive experiments conducted on three prominent datasets unequivocally demonstrate the superior performance of our framework. Our innovative approach yields a noteworthy enhancement, manifesting as an average reduction of 0.5° (equivalent to a substantial 9.3% improvement) in angular error.
@@ 2.1 Gaze Estimation Evolution
Early approaches attempt to achieve gaze estimation by predicting a mass of points on a 2D screen[55]. Owing to significant variations in the positions of cameras across different devices, the generalization capability of the 2D gaze estimation model is notably constrained. To mitigate this issue, a series of 3D gaze estimation approaches[11, 12, 13][17][23, 24, 25] are proposed to infer person gaze in the real world by enriching geometric information, such as varying shooting conditions. In contrast to geometric-based approaches, appearance-based models have garnered escalating attention within the research community. This is attributed to the capability of gaze regressors of this type to utilize images captured by conventional cameras as input. Benefiting from the huge success of deep learning[45][56], the performance of gaze estimation algorithms has been substantially improved [22]. These methods establish mapping functions between raw images and gazes through the powerful nonlinear fitting ability of neural network, and can achieve high accuracy even on unconstrained settings.
Specifically, the approach [22] unprecedentedly leverages high-capacity CNN to regress gazes from the facial image inputs. Zhang et al. [17] introduce a spatial attention mechanism to emphasize on the facial areas within input scenes, effectively suppressing the noises from redundant image regions. Meanwhile, Cheng et al. [23] realize the asymmetry between two eyes and further propose an asymmetric regression backbone consisting of four CNN stems. Stimulated by atrous convolution in image classification[57], Chen et al. [24] strive to enlarge receptive fields through dilated convolutions without incurring extra computational overheads. Wang et al[25] present a unified framework including both adversarial learning and a Bayesian approach, thereby enhancing the transferability and accuracy of gaze estimation models. The CA-Net [11] delves into the coarse-to-fine estimation procedure. In particular, the coarse stage hammers at deducing a fundamental gaze direction from a facial image, whereas the subsequent step further refines the predicted direction based on the provided eye images. More recently, [51] introduce an unsupervised domain adaptation approach to estimate gaze that is not restricted by source limitations which is superior in cross-domain gaze estimation methods. In [52], cameras with diverse viewpoints are utilised to capture the subject’s image. Following that, they set out a dual viewpoint gaze estimation network. Albeit achieving impressive results, all these gaze estimators ignore the promising semantics merits provided by large-scale modern language models (e.g., BERT [32] and CLIP [18]). Therefore, there is a large room for performance improvement in gaze estimation.
@@ 2.2 Language-steering Visual Models
CLIP is trained through contrastive learning to maximize the cosine similarity between text embedding and image embedding of positive sample pairs. Such a training method breaks the long-standing heavy reliance on image labels in the visual field. The pre-trained CLIP models can chieve superior performance compared to those fully supervised methods in many classic visual tasks even in few-shot or zero-shot settings. Given the powerful transferable ability of CLIP, many studies have explored how to apply CLIP to various specific downstream tasks. Several methods [20][26, 27, 28, 29]have achieved promising performance with the help of CLIP model on 3d avatar generation, age estimation, image aesthetics assessment, semantic segmentation and so on. These methods usually apply pre-trained CLIP model as backbone, and design a language-image interaction mechanism based on specific tasks after defining different prompts which usually are general language expression of the target images in advance.
@@ 3 Methodology
In this section, we describe how we perform gaze estimation with text knowledge. Firstly, we will illustrate the CLIP model which contains rich semantic knowledge of images and texts through unsupervised pre-training. Next, we will introduce the details in GazeCLIP.
@@ 3.1 Revise for CLIP
Influenced by large models in the field of natural language processing[30, 31, 32, 33], some research has begun to study how to apply the ideas of pre-training and large models to the field of computer vision, and break away from the dependence on large amounts of label data during visual task training since unsupervised learning models will have stronger transfer capability typically. CLIP[18], which collects 400 million pairs of image and text on the internet for model pre-training, has established the connection between visual concept and language concept. All of these texts are natural language descriptions of images. Specifically, during pre-training, given a batch of image-text pairs, they are fed into the image encoder and text encoder respectively, and then the cosine similarity of the positive sample pair is maximized through contrastive learning. By mapping the features of two modalities to a joint embedding space, the pre-trained CLIP model owns strong transfer ability, enabling it to be applied well to various downstream visual tasks even in zero-shot setting. For example, as image classification, a series of class labels (e.g., ‘car,’ ‘plane’) can be constructed to a prompt template rather than a single word since a complete sentence description is used in CLIP pre-training. The most common example is “a photo of [CLASS]”. The prompt is then fed into the text encoder to create class embeddings, which are utilized for computing similarity with image embeddings for classification. When applying the CLIP model to other downstream tasks, two issues need to be considered: firstly, constructing appropriate prompt that match the image, and secondly, effectively integrating the image embeddings and text embeddings. In the original CLIP model, the focus is on the similarity between the image and the text which is still a classification task essentially. Therefore, it needs to be adjusted for different tasks. In this work, we are the first to explore how to transfer the rich image and text knowledge learned by CLIP to the task of gaze estimation.
Figure 2: Architecture of the GazeCLIP. We only use facial images as input and take image encoder and text encoder in CLIP as backbone to extract feature. During training, we fine-tune the image encoder while the parameters of the text encoder are frozen.

### v3

@@ Abstract
Over the past decade, visual gaze estimation has garnered increasing attention within the research community, owing to its wide-ranging application scenarios. While existing estimation approaches have achieved remarkable success in enhancing prediction accuracy, they primarily infer gaze from single-image signals, neglecting the potential benefits of the currently dominant text guidance. Notably, visual-language collaboration has been extensively explored across various visual tasks, such as image synthesis and manipulation, leveraging the remarkable transferability of large-scale Contrastive Language-Image Pre-training (CLIP) model. Nevertheless, existing gaze estimation approaches overlook the rich semantic cues conveyed by linguistic signals and the priors embedded in CLIP feature space, thereby yielding performance setbacks. To address this gap, we delve deeply into the text-eye collaboration protocol and introduce a novel gaze estimation framework, named GazeCLIP. Specifically, we intricately design a linguistic description generator to produce text signals with coarse directional cues. Additionally, a CLIP-based backbone that excels in characterizing text-eye pairs for gaze estimation is presented. This is followed by the implementation of a fine-grained multi-modal fusion module aimed at modeling the interrelationships between heterogeneous inputs. Extensive experiments on three challenging datasets demonstrate the superiority of the proposed GazeCLIP which achieves the state-of-the-art accuracy.
@@ 1 Introduction
During the past decade, visual gaze estimation has garnered ever-increasing attention thanks to its widespread application scenarios, such as saliency detection [47], virtual reality[51], human-robot interaction[19, 44], medical diagnosis[4], driver fatigue estimation[53] and other areas[37, 13]. Gaze estimation aims to accurately infer the direction of a person’s gaze from an input image, typically focusing on the human face or eyes. Despite its promising practical potential applications, this visual task faces significant challenges. For instance, the utilization of diverse camera angles inevitably leads to unpredictable variations in lighting, color, and facial features, which can severely degrade performance. To struggle with this issue, several existing studies[59, 7, 5, 46, 3, 6, 1, 48] have made lots of efforts to marry the stability and robustness to regression algorithms. An additional catalyst for the thriving of investigations on gaze estimation is the increasing availability of large-scale datasets [15, 21, 16, 61, 58] which consist of varied subjects including lighting conditions, head poses and so on.
@@
Inspired by the huge success of deep learning in computer vision[24, 18, 40], an increasing number of researchers are shifting their focus towards modeling intricate mappings between input signals and gaze labels[58, 59, 57]. Nonetheless, these methods typically operate on single-modal facial images as input, making them susceptible to the challenge of isolating gaze-specific features amidst unrelated regions (e.g., the nose and mouth) scattered throughout the input image as different regions are prone to confuse the model on what to prioritize in learning. A general model for remote POG estimation based on video[17] evades this problem by mining geometric characteristics of the eye. Albeit the impressive progress, conventional geometry-modelling approaches exhibit several limitations such as reliance on controlled laboratory environments and poor generalization performance to new subject. This has led to a surge in the booming of appearance-based gaze estimation algorithms, harnessing the capabilities of cutting-edge deep learning models. To enhance the estimation performance, recent learning-based protocols [9, 31, 8] excel at extracting useful features from raw facial or eye images through the design of diverse high-capacity CNN models, see Figure 1(a). Furthermore, the inevitable drastic variations introduced by diverse camera views and individual characteristics (e.g. gender, skin color, and glasses) [15, 57] can easily lead the model to fail in learning highly non-nonlinear relationships, resulting in overfitting and sub-par performance. Therefore, there is a large room for further exploration of gaze regression model with enhanced learning capabilities.
@@
Recently, a cohort of large-scale pre-traiend linguistic-vision models [34, 12, 36, 14] have emerged as dominant forces in artificial intelligence, sparking a surge in multi-modal learning. Particularly, Contrastive Language-Image Pre-Training (CLIP) [35] elicits the foremost attention within the research community and has been widely adopted for various downstream vision tasks, such as image generation and visual question answering [38, 39, 41]. Specifically, CLIP’s backbone is comprised of a set of transformer blocks and is trained using an extensive collection of image-text pairs (400 million). The key advantage of incorporating CLIP lies in its implicit enrichment of features with the nuanced semantics embedded in natural language. For example, object detection [42], semantic segmentation [50], and image classification [62] have achieved notable success by harnessing the guiding capabilities of the prominent CLIP model.
@@
Despite the remarkable progress made in previous approaches, the field of gaze estimation has yet to fully exploit the substantial benefits of leveraging language guidance. Indeed, the task of regressing raw face/eye images into continuous gaze vectors presents an intractable challenge. Therefore, a "divide-and-conquer" approach naturally emerges, aiming to initially identify coarse-grained gaze directions and then utilize the coarse-granularity pseudo labels to guide fine-grained gaze estimation. Moreover, considering the potential decrease in estimation accuracy when the target gaze direction deviates from the camera’s central axis compared to direct gaze towards the camera, we posit that the estimation accuracy can be notably enhanced through the guidance of coarse-grained pseudo-directions. These coarse directions, represented as natural language prompts, can be encoded through pre-trained CLIP encoders, thereby enhancing the discriminative embeddings with robust guidance ability.
@@
While linguistic prompts effectively convey high-level and coarse-grained directions, devising appropriate prompts from facial/eye images poses a significant challenge, as articulating the nuances between input images in words is difficult. Unlike image classification tasks, the generation of meaningful language descriptions (such as gaze direction prompt) using the full-fledged BLIP [26] is inherently impractical as the training of BLIP fully ignores gaze face samples, thereby impeding the applicability of CLIP in gaze estimation. Due to the particularity of the gaze estimation dataset, natural languages generated using BLIP will always be "A man’s/woman’s face", but what we need for our approach is a rough direction guide, and it will not appear in common language descriptions.
@@
To address the aforementioned challenges, this paper delves into the intricate process of formulating suitable prompts and proceeds to carefully design a language-vision protocol (GazeCLIP). This protocol aims at enhancing the semantic space of the gaze estimation model. Specifically, we pre-define the text prompt as “A photo of a face gazing [class]”, where [class] represents a set of base direction(front, up, down, left, right). To make the constructed prompt conform to the language prior knowledge of the CLIP model, we directly use the zero-shot CLIP model to determine a direction for each image. We calculate the similarity between each input image and the sentence composed of the five directions mentioned above, and take the direction with the highest similarity as the text input. The main reason for this is that we hope to build a natural language that conforms to prior knowledge of the CLIP model. On top of the generated prompt-image pairs, we present a cross-attention-based module to align the representations from two heterogeneous modalities. Following the fine-grained fusion of multi-modal features through transformer blocks and cross-attention condensers, the resulting refined image embedding are combined with the original image embedding together to act as the input of final regression head. As depicted in Fig. 1, Unprecedentedly, our emphasis lies on harnessing the guidance provided by textual signals, rather than solely struggling to design image encoders.
@@
We evaluated our proposed model on three challenging public dataset benchmarks for gaze estimation. Extensive experiments demonstrate that GazeCLIP achieves state-of-the-art performance in all three datasets. Particularly, compared with existing approaches, the GazeCLIP gains significant improvement with an angular error decrease of 0.4° on average(4.1° → 3.6°, 7.4° → 7.3°, 5.3° → 4.7°). In a nutshell, the contributions of our GazeCLIP are threefold:
@@
  • In Use. A novel text-guided gaze estimation method is proposed in this paper, namely GazeCLIP, to enhance the performance of gaze estimation, which is inspired by strong generalization capability of CLIP model in various downstream vision tasks. To the best of our knowledge, this is the first attempt to distil rich knowledge from full-fledged pre-trained language-vision model for guiding the learning of gaze estimation network.
@@
  • SOTA Performance. Extensive experiments conducted on three prominent datasets unequivocally demonstrate the superior performance of our framework. Our innovative approach yields a noteworthy enhancement, manifesting as an average reduction of 0.4° (equivalent to a substantial 8% improvement) in angular error.
@@ 2.1 Gaze Estimation Evolution
Early approaches attempt to achieve gaze estimation by predicting a mass of points on a 2D screen[23]. Due to significant variations in the positions of cameras across different devices, the generalization capability of the 2D gaze estimation model is notably constrained. To mitigate this issue, a series of 3D gaze estimation approaches[8, 15, 21, 59, 7, 5, 46] are proposed to infer person gaze in the real world by enriching geometric information, such as varying shooting conditions. In contrast to geometric-based approaches, appearance-based models have garnered escalating attention within the research community. This is attributed to the capability of gaze regressors of this type to utilize images captured by conventional cameras as input. Taking advantage of the great success of deep learning[18, 43], the performance of gaze estimation algorithms has improved substantially.
Specifically, the approach [57] unprecedentedly leverages high-capacity CNN to regress gazes from the facial image inputs. Zhang et al. [59] introduce a spatial attention mechanism to emphasize facial areas within input scenes, effectively suppressing noise from redundant image regions. Meanwhile, Cheng et al. [7] realize the asymmetry between two eyes and further propose an asymmetric regression backbone consisting of four CNN stems. Stimulated by atrous convolution in image classification[54], Chen et al. [5] strives to enlarge receptive fields through dilated convolutions without incurring extra computational overhead. Wang et al[46] present a unified framework including both adversarial learning and a Bayesian approach, thereby enhancing the transferability and accuracy of gaze estimation models. The CA-Net [8] delves into the coarse-to-fine estimation procedure. In particular, the coarse stage hammers at deducing a fundamental gaze direction from a facial image, whereas the subsequent step further refines the predicted direction based on the provided eye images. [2] began to introduce an attention mechanism to extract key features of eye images. More recently, [3] introduced an unsupervised domain adaptation approach to estimate gaze that is not restricted by source limitations, which is superior in cross-domain gaze estimation methods. In [6], cameras with various viewpoints are used to capture the subject’s image. Following that, they set out a dual viewpoint gaze estimation network. [52, 32, 11, 29, 30] are also concerned about the performance under cross-dataset settings and aim to build more generalized models for gaze estimation task. [49] has demonstrated that higher resolution images will have a positive impact on the results, even with a simple backbone network. Albeit achieving impressive results, all these gaze estimators ignore the promising semantics merits provided by large-scale modern language models (e.g., BERT [12] and CLIP [35]). Therefore, there is a large room for performance improvement in gaze estimation.
@@ 2.2 Language-steering Visual Models
CLIP is trained through contrastive learning to maximize the cosine similarity between text embedding and image embedding of positive sample pairs. Such a training method breaks the long-standing heavy reliance on image labels in the visual field. The pre-trained CLIP models can chieve superior performance compared to those fully supervised methods in many classic visual tasks even in few-shot or zero-shot settings. Given the powerful transferable ability of CLIP, many studies have explored how to apply CLIP to various specific downstream tasks. Several methods [50, 20, 27, 28, 55]have achieved promising performance with the help of the CLIP model on 3d avatar generation, age estimation, image aesthetics assessment, semantic segmentation, and so on. These methods usually apply pre-trained CLIP model as backbone, and design a language-image interaction mechanism based on specific tasks after defining different prompts which usually are general language expression of the target images in advance.
@@ 3 Methodology
In this section, we outline our approach for performing gaze estimation with text knowledge. We begin by illustrating the CLIP model, which contains rich semantic knowledge of images and texts through unsupervised pre-training. Next, we delve into the details in GazeCLIP.
@@ 3.1 Revise for CLIP
Influenced by the success of large models in the field of natural language processing[33, 34, 12, 36], research has begun to study how to apply the ideas of pre-training and large models to the field of computer vision. This shift aims to reduce dependence on large amounts of labeled data during visual task training, as unsupervised learning models typically exhibit stronger transfer capabilities. CLIP[35], exemplifies this approach, leveraging 400 million pairs of image-text data collected from the internet for pre-training. These texts serve as natural language descriptions of images. During pre-training, a batch of image-text pairs is fed into separate image and text encoders, and the contrastive learning framework maximizes the cosine similarity of positive sample pairs. By mapping features of two modalities to a joint embedding space, the pre-trained CLIP model possesses strong transferability, making it applicable to various downstream visual tasks, even in zero-shot settings. For instance, in image classification, a series of class labels (e.g., ‘car’, ‘plane’) can be constructed to a prompt template, such as “a photo of [CLASS]”, as utilized in CLIP pre-training. This prompt is then fed into the text encoder to create class embeddings, which are used to compute similarity with image embeddings for classification. When applying the CLIP model to other downstream tasks, two considerations arise: constructing appropriate prompts matching the image and effectively integrating image embeddings and text embeddings. While the original CLIP model focuses on image-text similarity, often for classification tasks, adjustments are necessary for different tasks. In this work, we are the first to explore transferring the rich image and text knowledge learned by CLIP to the task of gaze estimation.
Figure 2: Architecture of the GazeCLIP. We only use facial images as input and take the image encoder and text encoder in CLIP as the backbone to extract feature. During training, we fine-tune the image encoder while the parameters of the text encoder are frozen.